OpenStudy (anonymous):

Just a stupid question probably, but... Using Heron of Alexandria's algorithm to find the Square root of a number; how well does it work trying to find the square root X, if X=35? I find it to be flawed in this case. Am I just over thinking it?

OpenStudy (anonymous):

The algorithm creates a never ending cycle of wrong answers in this case IMO. I'm sure that 35 wouldn't be the only instance of this failure.

OpenStudy (anonymous):

6 The answer using this algorithm is obviously going to create an answer of 36 which is greater than 35

OpenStudy (anonymous):

I guess the answer I am looking for is this, is 6 close enough? Actual answer would obviously be 5.9160797831

OpenStudy (anonymous):

I think there may be other flaws on some, for example if you carry on after 'getting close' you can generate a complete pile of mumbo jumbo !

OpenStudy (anonymous):

Okay so it was a stupid question LOL. I just didn't double check my own results. I Started by looking for the SqR of 35 using the algorithm. I guessed by starting with G=5 , G*G=25, X/G=7, 1/2(G+X/G)=6, Repeat G=6, G*G=36, X/G=5.83333333333333 which is pretty close to the correct answer.